OpenStudy (anonymous):

Solve the following quadratic equation using the quadratic formula and then choose the correct solution set. 6x^2 - 7x + 2 = 0

Directrix (directrix):

First, we need to know the QF (quadratic formula).

OpenStudy (anonymous):

That is \[X=-b \pm \sqrt{b ^{2}-4ac}\]

Directrix (directrix):

That is correct if you place it over (2a). I have the formula attached. For 6x^2 - 7x + 2 = 0. what are the values of a, b, and c?

OpenStudy (anonymous):

A=6 B=7 C=2

Directrix (directrix):

Attached is how I set up QF problems to avoid sign problems.

Directrix (directrix):

A=6 B=7 -----------> This should be -7 C=2

OpenStudy (anonymous):

so its set up like this \[-7\pm \sqrt{-7-4*6*2}\]

Directrix (directrix):

Here's what I have so far.

OpenStudy (anonymous):

its that right

Directrix (directrix):

I don't think so.

OpenStudy (anonymous):

ya me too

Directrix (directrix):

Under the radical, the b has to be squared: (-7)^2 = 49

Directrix (directrix):

I am writing the substitutions for a, b, and c, on the attached files. Have you seen my work?

OpenStudy (anonymous):

ya

Directrix (directrix):

Look at this:

Directrix (directrix):

I think your formula was off a bit. Then, the substitution for a, b, and c was off. Writing all the steps takes time but I recommend it if you are having trouble with the problem.

Directrix (directrix):

@Shannon20150 We left off with x = 8/12 and x = 6/12. Those values will simplify to what?

Directrix (directrix):

8/12 = 2/? and 6/12 = 1/?